OpenStudy (anonymous):

solve the equation by using quadratic formula. 2x^2+12x+14=0

OpenStudy (anonymous):

the quadratic formula is \[x= \frac{ -b \pm \sqrt{b ^{2}}-4ac }{ 2a }\]

OpenStudy (anonymous):

so find a,b, and c in the equation. a=2, b=12, and c=14 then you just have to plug them into the formula like this: \[\frac{ -12\pm \sqrt{12^{2}}-4(2)(14) }{ 2(2) }\] then you solve the equation for x

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\]

OpenStudy (espex):

As @winterfrez pointed out, the b^2-4ac is all under the radical so your equation would be: \[\frac{ -12\pm \sqrt{{12^{2}}-4(2)(14)} }{ 2(2) }\]